OpenStudy (anonymous):

The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (alexandervonhumboldt2):

pyth theorem is a^2+b^2=c^2

OpenStudy (alexandervonhumboldt2):

6^2+8^2=10^2 check this 5^2+12^2=13^2 check this too

OpenStudy (anonymous):

6^2 = 36 8^2 = 64 10^2 = 100 ==> Reverse pythagorean theorem 36 + 64= 100 so these are pythagorean triples same goes for the next ones

OpenStudy (anonymous):

A Pythagorean Triple is a set of positive integers, a, b and c that fits the rule in this case you have 6, 8, and 10

OpenStudy (anonymous):

So what's the answer? I don't understand :(